OpenStudy (anonymous):

check my answer? Which of the following represents the set of integers greater than or equal to −5? {−5, −6, −7, …} {−5, −4, −3, …} {−5, −4, −3, …, 3, 4, 5} {all positive numbers}

OpenStudy (anonymous):

i got the first one is it right?

hartnn (hartnn):

no,integers greater than -5 are, -5+1,-5+2,-5+3......

OpenStudy (hba):

Yeah :)

OpenStudy (anonymous):

thank you(:

hartnn (hartnn):

hba ?

hartnn (hartnn):

the first one is incorrect.

OpenStudy (hba):

I mean that @sylinan you are wrong @hartnn is right.

OpenStudy (hba):

All the negative and positive numbers are included.

OpenStudy (hba):

Negative integers and positive integers.

OpenStudy (anonymous):

ohhh haha so it would be the third one?

hartnn (hartnn):

no....its -5+1=-4 -5+2=-3 means -4,-3,-2,....

hartnn (hartnn):

no, its the 2nd one.....

OpenStudy (anonymous):

@hartnn ohh okay thank you(:

hartnn (hartnn):

because the set will be infinite.

hartnn (hartnn):

ok, welcome ^_^